# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v3

@@ Abstract
The universal approximation theorem and foundational results in deep reinforcement learning establish that all human labor, cognitive and physical, is in principle automatable. The critical question is not whether full automation will arrive, but who will control it. This paper introduces anarcho-accelerationism: a sociotechnical framework for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. We propose the Liberation Stack, a layered architecture of energy, manufacturing, food, communication, knowledge, and governance commons, powered by frontier clean energy technologies within an accelerationist ecologism that achieves sustainability through technological abundance rather than degrowth. We demonstrate that anarcho-accelerationism builds bridges to liberalism, socialism, environmentalism, feminism, cooperativism, faith traditions, and the hacker ethic. We propose Universal Desired Resources (UDR) as a post-monetary design principle and show that UDR eliminates the material basis of all axes of structural inequality identified by intersectional theory. Drawing on Maslow’s hierarchy of needs, we show that the Liberation Stack satisfies basic human needs universally, enabling what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy. We propose a progressive transition from Universal Basic Income to UDR, leveraging capitalist automation while building commons alternatives. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, and guifi.net confirms that commons-based systems already operate at scale. We conclude that one of the most important engineering challenges of the 21st century is building the decentralized infrastructure through which robotic automation serves all of humanity, enabling a society centered on care, love, and the common good.
Keywords: anarcho-accelerationism; full automation; commons governance; universal desired resources; accelerationist ecologism
1 Introduction: The Convergence Window
@@
The dominant architectures of our sociotechnical systems (centralized cloud platforms, extractive business models, proprietary codebases, fossil-fuel energy grids, top-down governance) were designed to concentrate power. They are spectacularly effective at this task. But they are not the only architectures available to us. A growing body of evidence from computer science, institutional economics, cooperative studies, and political ecology suggests that decentralized, commons-based, federated architectures can achieve equal or superior performance in coordination, resilience, innovation, and human well-being (Benkler, 2006; Ostrom, 1990; Scholz, 2016).
@@
This paper introduces anarcho-accelerationism: a framework for sociotechnical system design that takes this evidence seriously. The name is deliberately provocative, but the substance is rigorous. The prefix anarcho- refers not to chaos but to the 150-year tradition of anarchist political philosophy, from Kropotkin’s empirical study of mutual aid in nature (Kropotkin, 1902) to Bookchin’s ecological municipalism (Bookchin, 2005) to Ostrom’s Nobel Prize-winning work on commons governance (Ostrom, 1990). The core principle is self-organization without domination. The suffix -accelerationism refers not to reckless technological enthusiasm but to the critical insight that current productive forces, if redirected from accumulation toward liberation, are sufficient to abolish artificial scarcity and enable universal flourishing (Srnicek and Williams, 2015). The core principle is technological sufficiency for post-scarcity.
The synthesis yields a design philosophy centered on three values that may sound naive to computer scientists trained in optimization and efficiency, but which we will show are both technically implementable and empirically validated: care, understood as the design of systems that sustain life, relationships, and communities rather than maximize engagement, extraction, or growth (Federici, 2012; Puig de la Bellacasa, 2017); love, understood not as sentimentality but as political commitment, the design choice to build systems that strengthen bonds between people rather than atomize them (hooks, 2000; Fromm, 1956); and the common good, the conviction that resources essential for human dignity, such as energy, food, knowledge, communication, and code, should be governed as commons, not enclosed as property (Ostrom, 1990; Hess and Ostrom, 2007).
This urgency is not abstract. As we demonstrate in Section˜2, foundational theorems in machine learning (the universal approximation theorem and the convergence results of deep reinforcement learning) guarantee that all human labor, cognitive and physical, is in principle automatable. The empirical evidence, from LLM-powered coding agents to humanoid robots, confirms that this automation is not a distant prospect but an ongoing process. The question is no longer whether jobs will disappear, but who will control the machines that replace them.
The central claim of this paper is that anarcho-accelerationism is not a partisan ideology but a bridge-building framework. It finds genuine common ground with every major political and intellectual tradition of our time (Section˜4), demonstrates that UDR achieves universal intersectional liberation by eliminating the material basis of all axes of oppression, satisfies the full hierarchy of human needs as described by Maslow, and provides a progressive transition path from the current system through synthetic liberty (Section˜5). It is grounded in a rigorous historical methodology drawn from the major historians of anarchism (Section˜3). It offers a technical architecture (Section˜6) powered by frontier clean energy technologies within an accelerationist ecological framework, with comprehensive scientific evidence for the ecological superiority of technological acceleration over degrowth (Section˜7), and grounded in empirical evidence of working commons (Section˜8). We believe this matters for computer science because we are the ones building the infrastructure. Every protocol we design, every platform we deploy, every algorithm we optimize embodies a political choice about who has power and who does not (Winner, 1980; Lessig, 2006). We can choose differently.
2 The Automation Imperative: Why This Framework, Why Now
The urgency of anarcho-accelerationism stems from a technological reality that is no longer speculative: full automation of human labor is theoretically possible in principle and empirically underway. We ground this claim in two foundational results from machine learning theory and their observable consequences.
@@ 2 The Automation Imperative: Why This Framework, Why Now
The universal approximation theorem, established independently by Cybenko (Cybenko, 1989) and Hornik et al. (Hornik, 1991), proves that a feedforward neural network with a single hidden layer containing a finite number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary precision. Formally, for any continuous function $f:K\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$ and any $\epsilon>0$, there exists a neural network $\hat{f}$ such that $\|f(x)-\hat{f}(x)\|<\epsilon$ for all $x\in K$. The implication is significant: every task that can be described as a continuous mapping from inputs to outputs can in principle be represented by a neural network. We note important caveats: the UAT is an existence result that guarantees an approximating network exists but provides no constructive procedure for finding it, no bound on the network size required, and no guarantee that gradient-based training will converge to the approximator in tractable time. The theorem applies only to continuous functions on compact subsets, and the number of training samples required (sample complexity) may be prohibitively large. Our claim is therefore that full automation is possible in principle: the theoretical barrier has been removed, not that it is trivially achievable for all tasks.
@@
The convergence theorems for deep reinforcement learning (DRL) (Sutton and Barto, 2018; Mnih et al., 2015) establish that an agent interacting with any environment describable as a Markov Decision Process (MDP) can learn an optimal policy through observation of states, actions, and rewards. For any MDP $(S,A,P,R,\gamma)$ with finite state and action spaces, Q-learning converges to the optimal action-value function $Q^{*}$ given sufficient exploration (Watkins and Dayan, 1992). With deep neural networks as function approximators (DQN and its successors), Q-learning has achieved remarkable empirical success in continuous, high-dimensional state spaces (Mnih et al., 2015), though we note that formal convergence guarantees do not extend from the tabular case to function approximation in general: the combination of bootstrapping, off-policy learning, and function approximation (the “deadly triad”) can cause divergence (Sutton and Barto, 2018). The practical successes summarized in Table˜1 are therefore empirical achievements, not deductive consequences of the tabular convergence theorem. Combined with imitation learning, where policies are acquired by observing human demonstrations rather than through explicit reward signals (Schaal, 1999), this suggests that any physical task that a human can perform could, in principle, be learned by a robotic agent that observes the human performing it, subject to caveats including partial observability of the demonstrator’s internal states, distributional shift from compounding errors, embodiment mismatch between human and robot kinematics, and the sample complexity of learning dexterous manipulation.
The progression of DRL milestones, summarized in Table˜1, illustrates how rapidly these theoretical guarantees are translating into practical capabilities. In little more than a decade, learning agents have gone from playing Atari games to defeating world champions in the most complex board games, solving fundamental problems in biology, and controlling humanoid robots in unstructured environments.
@@
2017 AlphaGo Zero (Silver et al., 2016) Superhuman Go from self-play alone, no human data, pure reinforcement learning
@@
2019 AlphaStar Grandmaster-level StarCraft II, real-time strategy with imperfect information
@@
These are not theoretical curiosities. Large language models (LLMs) have crossed a critical capability threshold. Systems such as GPT-4 (OpenAI, 2023) and Claude (Anthropic, 2024), together with their open-source counterparts, now write production-quality software through AI coding agents such as GitHub Copilot and Claude Code; draft legal contracts and summarize case law at a fraction of the cost of human attorneys; diagnose medical images with accuracy matching or exceeding specialists; generate financial analyses and detect fraud patterns across massive datasets; and assist in hypothesis generation, experimental design, and data analysis across every scientific discipline.
Multiple independent studies, summarized in Table˜2, converge on the same conclusion: the majority of the current labor force faces significant automation exposure. These estimates predate the most capable current models, suggesting that the actual impact will be substantially larger.
@@
Robotics, long the slower partner in the automation trajectory, is entering an exponential phase driven by the same deep learning breakthroughs. Humanoid robots such as Boston Dynamics Atlas, Tesla Optimus, and Figure 01 now perform warehouse logistics, manufacturing assembly, and unstructured physical tasks. Agricultural robots autonomously plant, monitor, and harvest crops. Autonomous vehicles navigate complex, unstructured traffic environments. Construction robots lay bricks, pour concrete, and perform structural inspection. Domestic robots clean, cook, and assist with elder care. Figure˜1 shows the automation risk by occupational sector, illustrating that physical labor sectors face some of the highest exposure, but that no sector is immune.
@@
Figure 1: Probability of automation by occupational category, adapted from Frey and Osborne (2017). Both blue-collar sectors (transportation, production, food service) and white-collar sectors (office, sales, legal) face substantial automation risk. No occupational category is fully immune.
The theoretical results (universal approximation and tabular DRL convergence) establish that no mathematical barrier prevents the automation of any task that can be formalized as a continuous function or finite MDP, and the empirical trajectory suggests that practical barriers are falling rapidly: every job, whether it involves manipulating symbols or manipulating matter, is automatable in principle and increasingly in practice.
@@
Figure 2: The automation fork. Full automation of labor is theoretically possible in principle and empirically underway. The outcome depends on the ownership and governance architecture: centralized ownership leads to neo-feudalism; decentralized commons governance leads to a society of universal care.
Path B is a care-centered society: a society where the fundamental organizing principle is not production for profit but care for persons. This is not a utopian fantasy; it is the logical consequence of decentralized full automation combined with commons governance. But the implications of full automation extend even further than the choice between centralized and decentralized ownership. We argue that full automation, combined with advances in molecular synthesis and recursive artificial intelligence, leads to consequences that fundamentally reshape the very categories through which we think about economics, ethics, and social organization.
The first consequence is the obsolescence of money. Money exists as a coordination mechanism for allocating scarce goods produced by human labor. When all production is automated and the marginal cost of most goods approaches zero (as it already has for digital goods), the function of money dissolves. Universal Basic Income (UBI), widely proposed as a solution to technological unemployment (Srnicek and Williams, 2015), remains trapped within the monetary paradigm: it redistributes purchasing power but preserves the market architecture that generates inequality in the first place. We propose instead the concept of Universal Desired Resources (UDR): in a fully automated commons, every person has direct access to the goods and services they desire, requested from the robotic productive commons, subject only to ecological sustainability constraints. UDR is not a transfer payment; it is a design principle for a post-monetary society. The difference is fundamental: UBI asks “how do we give people money when there are no jobs?”; UDR asks “why do we need money when robots can produce everything?”
The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The Liberation Stack (Section˜6) describes existing technologies, but the trajectory points further. Advances in additive manufacturing (3D printing at molecular scales), synthetic biology (DNA-programmable organisms and materials), and recursive AI self-improvement (where AI systems design better AI systems, better robots, better energy systems, and better versions of themselves) converge toward a productive infrastructure capable of synthesizing any physical object, any material, any energy configuration, bounded only by the laws of physics rather than the limits of human labor or capital. 3D bioprinters already fabricate living tissue; DNA origami techniques fold genetic material into nanoscale structures; large language models already write the code that trains the next generation of models. The logical endpoint is not merely automation of existing jobs but the creation of a universal fabrication commons: a decentralized network of robotic systems capable of producing, on demand, whatever any community or individual needs, from food to medicine to housing to energy infrastructure.
The third consequence is an ethical clarification: it is not only permissible but morally required to delegate all instrumental labor to machines, on the basis of a clear ontological distinction between computational intelligence and phenomenal consciousness. As argued in prior work, phenomenal consciousness (subjective experience, qualia, the capacity to suffer) is independent of computational intelligence (Garrido-Merchán and Lumbreras, 2022). On this view, machines, however sophisticated their information processing, do not experience phenomenal consciousness (Garrido-Merchán, 2024); they do not understand in the human sense of the term (Garrido-Merchán and Blanco, 2022). We acknowledge that this position, while philosophically grounded, engages with the hard problem of consciousness, which remains unresolved; functionalist or illusionist accounts of consciousness could in principle reach different conclusions. A robot that assembles furniture does not feel boredom. An AI that writes code does not feel alienation. A harvesting machine does not suffer exhaustion. Delegating labor to entities that cannot suffer, in order to liberate entities that can, is not exploitation; it is care applied at civilizational scale.
The fourth consequence is that morality itself becomes computationally tractable. If the ethical framework of the care-centered society is grounded in humanist values (dignity, solidarity, sustainability, freedom from suffering), then the question of what constitutes moral behavior in specific situations can be formalized as a reinforcement learning problem. As shown by Garrido-Merchán and Lumbreras-Sancho (2023), human moral decision-making can be modeled as a Markov Decision Process where the reward function encodes humanist values and the state space captures the social and environmental context. The resulting optimal policy, the moral policy, maximizes the common good subject to humanist constraints. This does not mean that morality is delegated to machines; it means that computational tools can help communities discover, refine, and implement ethical principles with the same rigor that they apply to engineering problems. The DRL framework provides a theoretical basis for studying how such a moral policy might be discovered through experience, though we acknowledge that this formalization requires strong assumptions: that a reward function faithfully encoding humanist values can be specified (the open problem of value alignment), that the morally relevant state is approximately Markovian, and that the moral environment is sufficiently stationary. We present this as a research direction for helping communities reason about ethical trade-offs, not as a solved problem.
Taken together, these four consequences yield a vision that goes well beyond conventional proposals for technology-assisted social reform. We are not proposing a better welfare state, a more humane capitalism, or a fairer distribution of existing wealth. We are proposing that the combination of decentralized full automation, commons governance, and the ontological impossibility of machine suffering makes possible, for the first time in history, a society in which no conscious being is compelled to perform instrumental labor, and every conscious being has access to whatever they need to live a life of care, creativity, and meaning. The remainder of this paper provides the political bridges (Section˜4), the technical architecture (Section˜6), and the empirical evidence (Section˜8) for building this society.
Anarcho-accelerationism rests on four intellectual pillars. Kropotkin’s Mutual Aid: A Factor of Evolution (Kropotkin, 1902) challenged the Social Darwinist misreading of natural selection by documenting extensive cooperation across species. Contemporary evolutionary biology has vindicated this view: multilevel selection theory (Wilson, 2007), reciprocal altruism (Trivers, 1971), and cultural group selection (Henrich, 2016) all demonstrate that cooperation is not an anomaly but a primary driver of complex social organization. In computer science, this principle is realized in commons-based peer production (Benkler, 2006): large-scale collaboration without property rights or hierarchical command. Linux, with over 20,000 contributors (Linux Foundation, 2023), and Wikipedia, with 300,000 active editors producing 60 million articles in 300 languages, demonstrate that mutual aid scales to planetary coordination.
Ostrom’s (1990) institutional analysis demolished Hardin’s “tragedy of the commons” by documenting thousands of cases where communities successfully manage shared resources without either privatization or state control. Her eight design principles for long-enduring commons (clear boundaries, proportional equivalence, collective-choice arrangements, monitoring, graduated sanctions, conflict-resolution mechanisms, minimal recognition of rights, and nested enterprises) provide a rigorous institutional framework that maps directly onto distributed system design (De Moor, 2010).
Following Stiegler (Stiegler, 2010), we adopt the concept of pharmakon: every technology is simultaneously remedy and poison. The internet connects and surveils. AI augments and displaces. Automation liberates and precaritizes. This is not a reason for technological pessimism; it is a reason for technological governance. The question is never whether to use technology but who controls it, for what purposes, and under what institutional arrangements (Feenberg, 2002).
Drawing from the anarchist tradition of prefigurative politics (Graeber, 2013), we propose that the systems we build should embody the values they aim to promote. A platform designed for democratic governance should itself be governed democratically. A tool for decentralization should not run on a centralized server. A system promoting care should be maintained with care. This is not idealism; it is consistency of design, a principle that any software engineer recognizes as essential to system integrity.
3 Historical and Methodological Foundations

### v5

@@ Abstract
Foundational results in machine learning establish that all human labor is in principle automatable. This paper introduces peaceful anarcho-accelerationism, a rigorously defined sociotechnical framework grounded in the 200-year anarchist tradition from Godwin through Kropotkin to Bookchin, and in the methodological categories of Eltzbacher, Nettlau, and Corrêa, for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. We state five formal hypotheses and six research objectives, present a formal definition through analytical categories of interdependent spheres, and propose the Liberation Stack as a layered technical architecture with explicit preconditions and gate conditions for each layer. We introduce Universal Desired Resources as a post-monetary design principle that eliminates the material basis of intersectional oppression, and develop a framework for progressive state dissolution through incremental, reversible commons-building compatible with existing democratic institutions. A nonviolent social mobilization strategy maps concrete peaceful methods to each stage of transition. We show that accelerationism and degrowth share anarchist pacifism as substrate and differ only along a Pareto-optimal technological frontier. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, guifi.net, the Fediverse, and contemporary commons initiatives confirms that commons-based systems already operate at scale. We conclude with a phased roadmap specifying explicit assumptions, hard constraints, gate conditions between phases, and detailed limitations.
Keywords: anarcho-accelerationism; full automation; commons governance; universal desired resources; accelerationist ecologism; nonviolent strategy; Fediverse
1 The Convergence Window and the Automation Imperative
@@
Before proceeding, we state the formal hypotheses that this paper investigates and the research objectives it pursues, so that the reader can evaluate the argument against explicit claims rather than vague aspirations.
The first hypothesis (H1) is that all human labor, both cognitive and physical, is in principle automatable. This claim rests on the universal approximation theorem (Cybenko, 1989; Hornik, 1991) and the convergence results of deep reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2015; Watkins and Dayan, 1992), which together establish that no mathematical barrier prevents the automation of any formalizable task or any learnable sensorimotor policy. We emphasize that this is an in-principle claim: the existence of approximating networks and convergent policies does not guarantee tractable training, nor does it address tasks whose specification may be inherently incomplete. The second hypothesis (H2) is that decentralized commons governance of automation produces better outcomes than centralized corporate ownership, measured along three dimensions: inequality (Gini coefficient and wealth distribution), human well-being (life satisfaction, health outcomes, educational attainment, and Maslow-level satisfaction), and ecological footprint (carbon emissions, resource throughput, and biodiversity impact). This hypothesis is testable through comparative analysis of existing commons-based systems and through pilot implementations of the Liberation Stack. The third hypothesis (H3) is that nonviolent social mobilization can achieve the commons transition without state seizure, drawing on the empirical findings of Chenoweth and Stephan (2011) that nonviolent campaigns succeed at twice the rate of violent ones, and on the strategic framework of Sharp (1973) that identifies 198 methods of nonviolent action applicable to the commons transition. The fourth hypothesis (H4) is that the apparent opposition between degrowth and accelerationism is a false binary: both positions share anarchist pacifism as their ethical substrate and occupy different points along a Pareto-optimal frontier trading off ecological risk against human suffering during the transition. A mature movement would support both strategies simultaneously, and technological progress shifts the entire frontier toward the origin. The fifth hypothesis (H5) is that each layer of the Liberation Stack has identifiable preconditions for successful implementation, and that the layers form a partial ordering of dependencies: energy is the precondition for all material layers, communication is the precondition for knowledge and governance, and no layer can be sustainably deployed without the layers below it in the dependency graph.
The research objectives follow from these hypotheses. The first objective (O1) is to define peaceful anarcho-accelerationism rigorously through the analytical categories of Corrêa (Corrêa, 2015) and the comparative method of Eltzbacher (Eltzbacher, 1908), situating it within the 200-year anarchist tradition and demonstrating its continuity with classical principles. The second objective (O2) is to propose a concrete technical architecture, the Liberation Stack, with explicit preconditions and gate conditions for each layer, a dependency graph specifying the partial ordering of implementation, and a framework for progressive state dissolution through incremental commons-building. The third objective (O3) is to develop a nonviolent mobilization strategy grounded in the empirical evidence of Chenoweth and Stephan (2011) and the strategic repertoire of Sharp (1973), mapping specific nonviolent methods to each stage of the commons transition. The fourth objective (O4) is to demonstrate that intersectional liberation, as theorized by Crenshaw (1989) and Collins (2019), is a structural consequence of Universal Desired Resources: when the material basis of all axes of oppression is eliminated simultaneously, intersectional liberation follows as a design outcome rather than a political aspiration. The fifth objective (O5) is to provide empirical evidence from existing commons-based systems, including Linux, Wikipedia, Mondragon, Rojava, guifi.net, the Fediverse, and contemporary initiatives such as the Kudumbashree network, the Preston Model, Cooperation Jackson, and the Evergreen Cooperatives, demonstrating that the organizational principles of the framework are not speculative but operational. The sixth objective (O6) is to present a phased roadmap with explicit assumptions, preconditions, gate conditions between phases, hard constraints, detailed limitations, and directions for future research, so that the framework can be evaluated, tested, and refined by the broader scholarly community.
@@
The dominant architectures of our sociotechnical systems, namely centralized cloud platforms, extractive business models, proprietary codebases, fossil-fuel energy grids, and top-down governance, were designed to concentrate power. They are spectacularly effective at this task. But they are not the only architectures available to us. A growing body of evidence from computer science, institutional economics, cooperative studies, and political ecology suggests that decentralized, commons-based, federated architectures can achieve equal or superior performance in coordination, resilience, innovation, and human well-being (Benkler, 2006; Ostrom, 1990; Scholz, 2016).
@@
This paper introduces anarcho-accelerationism: a framework for sociotechnical system design that takes this evidence seriously. The name is deliberately provocative, but the substance is rigorous. The prefix anarcho- refers not to chaos but to the 150-year tradition of anarchist political philosophy, from Kropotkin’s empirical study of mutual aid in nature (Kropotkin, 1902) to Bookchin’s ecological municipalism (Bookchin, 2005) to Ostrom’s Nobel Prize-winning work on commons governance (Ostrom, 1990). The core principle is self-organization without domination. The suffix -accelerationism refers not to reckless technological enthusiasm but to the critical insight that current productive forces, if redirected from accumulation toward liberation, are sufficient to abolish artificial scarcity and enable universal flourishing (Srnicek and Williams, 2015). The core principle is technological sufficiency for post-scarcity. The synthesis yields a design philosophy centered on three values: care, understood as the design of systems that sustain life, relationships, and communities rather than maximize engagement, extraction, or growth (Federici, 2012; Puig de la Bellacasa, 2017); love, understood not as sentimentality but as political commitment, the design choice to build systems that strengthen bonds between people rather than atomize them (hooks, 2000; Fromm, 1956); and the common good, the conviction that resources essential for human dignity, such as energy, food, knowledge, communication, and code, should be governed as commons, not enclosed as property (Ostrom, 1990; Hess and Ostrom, 2007). As Nettlau wrote in surveying the entire history of anarchism, the movement is unified by a belief in reason and progress, and by the pursuit of the maximum possible realization of the liberty and wellbeing of all (Nettlau, 1996).
@@
The urgency of this framework stems from a technological reality that is no longer speculative: full automation of human labor is theoretically possible in principle and empirically underway. The universal approximation theorem, established independently by Cybenko (Cybenko, 1989) and Hornik et al. (Hornik, 1991), proves that a feedforward neural network with a single hidden layer containing a finite number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary precision. Formally, for any continuous function $f:K\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$ and any $\epsilon>0$, there exists a neural network $\hat{f}$ such that $\|f(x)-\hat{f}(x)\|<\epsilon$ for all $x\in K$. We note important caveats: the UAT is an existence result that guarantees an approximating network exists but provides no constructive procedure for finding it, no bound on the network size required, and no guarantee that gradient-based training will converge to the approximator in tractable time. Our claim is therefore that full automation is possible in principle: the theoretical barrier has been removed, not that it is trivially achievable for all tasks.
@@
The convergence theorems for deep reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2015) establish that an agent interacting with any environment describable as a Markov Decision Process can learn an optimal policy through observation of states, actions, and rewards. For any MDP $(S,A,P,R,\gamma)$ with finite state and action spaces, Q-learning converges to the optimal action-value function $Q^{*}$ given sufficient exploration (Watkins and Dayan, 1992). With deep neural networks as function approximators, Q-learning has achieved remarkable empirical success in continuous, high-dimensional state spaces, though formal convergence guarantees do not extend from the tabular case to function approximation in general. Combined with imitation learning, where policies are acquired by observing human demonstrations (Schaal, 1999), this suggests that any physical task that a human can perform could, in principle, be learned by a robotic agent, subject to caveats including partial observability, distributional shift, embodiment mismatch, and sample complexity. The progression of DRL milestones, summarized in Table˜1, illustrates how rapidly these theoretical guarantees are translating into practical capabilities.
@@
2017 AlphaGo Zero (Silver et al., 2016) Superhuman Go from self-play alone, no human data
@@
Large language models have crossed a critical capability threshold. Systems such as GPT-4 (OpenAI, 2023) and Claude (Anthropic, 2024), together with their open-source counterparts, now write production-quality software, draft legal contracts, diagnose medical images with accuracy matching specialists, generate financial analyses, and assist in hypothesis generation across every scientific discipline. Multiple independent studies converge on the same conclusion: the majority of the current labor force faces significant automation exposure, as summarized in Table˜2.
@@
Robotics is entering an exponential phase driven by the same deep learning breakthroughs. Humanoid robots such as Boston Dynamics Atlas, Tesla Optimus, and Figure 01 now perform warehouse logistics, manufacturing assembly, and unstructured physical tasks. Figure˜1 shows the automation risk by occupational sector.
@@
Figure 1: Probability of automation by occupational category, adapted from Frey and Osborne (2017). Both blue-collar and white-collar sectors face substantial automation risk. No occupational category is fully immune.
@@
Figure 2: The automation fork. Full automation of labor is theoretically possible in principle and empirically underway. The outcome depends on the ownership and governance architecture.
Path B is a care-centered society: a society where the fundamental organizing principle is not production for profit but care for persons. The implications of full automation extend further than this binary choice. The first consequence is the obsolescence of money. Money exists as a coordination mechanism for allocating scarce goods produced by human labor. When all production is automated and the marginal cost of most goods approaches zero, the function of money dissolves. We propose the concept of Universal Desired Resources (UDR): in a fully automated commons, every person has direct access to the goods and services they desire, requested from the robotic productive commons, subject only to ecological sustainability constraints. The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The third consequence is an ethical clarification: it is not only permissible but morally required to delegate all instrumental labor to machines, on the basis of a clear ontological distinction between computational intelligence and phenomenal consciousness (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022). A robot that assembles furniture does not feel boredom. An AI that writes code does not feel alienation. Delegating labor to entities that cannot suffer, in order to liberate entities that can, is care applied at civilizational scale. The fourth consequence is that morality itself becomes computationally tractable: human moral decision-making can be modeled as a Markov Decision Process where the reward function encodes humanist values (Garrido-Merchán and Lumbreras-Sancho, 2023), though this formalization requires strong assumptions about value alignment and moral stationarity.
Previous attempts at decentralized, commons-based social organization illuminate both the promise and the perils of the present moment. The Paris Commune of 1871 demonstrated that workers could govern a major city through direct democracy, establishing free education, gender equality in wages, and cooperative workshops, but it lasted only 72 days before being crushed by military force (Ross, 2015). The Makhnovshchina in Ukraine from 1918 to 1921 organized agricultural communes across a territory of seven million people under conditions of civil war, demonstrating the viability of federated anarchist governance at regional scale, but was destroyed by Bolshevik betrayal after having resisted both the White Army and foreign intervention (Malet, 1982). The Kronstadt rebellion of 1921 saw the very sailors who had enabled the Russian Revolution demand the soviet democracy they had been promised, only to be massacred by the regime they had helped create (Avrich, 1970). The Spanish Revolution of 1936 achieved the most extensive implementation of anarchist principles in history: the CNT-FAI collectivized factories, farms, and public services across Catalonia and Aragon, with industrial production maintained and in some cases increased under worker self-management (Bolloten, 1991). Agricultural collectives in Aragon organized production, distribution, and social services for hundreds of thousands of people through assembly-based governance (Casanova, 2006). Yet the revolution was destroyed by the convergence of Francoist military assault, Stalinist repression from within the Republican camp, and the CNT’s fateful decision to join the Republican government, contradicting its own anti-statist principles.
These movements failed not because their organizational principles were flawed, but because they lacked three conditions that are now, for the first time in history, within reach. First, they had no technological infrastructure for decentralized coordination at scale: no internet, no mesh networks, no federated protocols, no real-time communication across territories. The CNT’s extraordinary organizational achievement of reaching 1.5 million members relied on physical meetings, printed newspapers, and postal correspondence, all vulnerable to disruption by military force. Today, federated communication protocols (ActivityPub, Matrix), community mesh networks (guifi.net), and open-source governance tools (Decidim, Loomio) provide the infrastructure for coordination that the historical movements could only dream of. Second, they had no energy independence. Every historical anarchist territory depended on external energy sources controlled by hostile states. Community-owned solar microgrids, enhanced geothermal systems, and iron-air batteries now make energy sovereignty technically and economically feasible at the community scale. Third, they had no productive automation. The collectives of 1936 still required human labor for every productive task, making them vulnerable to the economic disruptions of war. Autonomous robotic systems for agriculture, manufacturing, and logistics now make economic self-sufficiency feasible without requiring the full participation of every community member in manual labor.
The remainder of this paper provides the formal definition (Section˜2), the political bridges (Section˜3), the technical architecture and ecological framework (Section˜4), the nonviolent mobilization strategy (Section˜5), the empirical evidence (Section˜6), and the roadmap (Section˜7) for building this society.
2 Theoretical Foundations: A Formal Definition of Peaceful Anarcho-Accelerationism